OpenStudy (anonymous):

How many hours are in a fortnight (2weeks)?

OpenStudy (anonymous):

336 Hours

OpenStudy (calculusxy):

One day = 24 hours One week = 7 days 2 weeks = (7*2) Hours = 24(7*2)